Tom Ilmanen

Address  Department der Mathematik, ETH Zentrum, CH-8092 Zürich, Switzerland.
Phone  +41-1-632-5443 (voice), +41-1-632-1570 (fax)
Email  c/o Hedi Oehler, oehler@math.ethz.ch
Mathematical  Geometric Analysis & Partial Differential Equations

Interests

Education  Haverford College, Pennsylvania, B.A. in Mathematics, 1984.

University of California at Berkeley, Ph.D. in Mathematics, 1991.

Employment  Visiting Member, Institute for Advanced Study, 1991 - 1992.

Visiting Assistant Professor, University of Wisconsin, Madison, 1992 - 1995. Assistant Professor, Northwestern University, 1995 - 1998. Visiting Scientist, Max-Planck Institut, Leipzig, 1997 - 1998. Ordinary Professor, Eidgenössiche Technische Hochschule, 1999 - present. Eilenberg Visiting Professor, Columbia University, Fall 2003. Eilenberg Visiting Professor, Columbia University, Spring 2005.

Thesis  “Elliptic Regularization and Partial Regularity for Motion by Mean Curvature.”
Advisor  Prof. L. C. Evans.
Lecture Series  Padova, 1994; ICTP, Trieste, 1995; Bochum, 1996; Keio, 1997; Freiburg, 1998;

Oberwolfach, 2000, Columbia 2003, 2005.

Honors & Fellowships Ranked twelfth in Putnam Exam, 1982.

NSF Postdoctoral Fellowship, 1993 - 1997.

Sloan Fellowship, 1996 - 2000.